# Full-scale Representation Guided Network for Retinal Vessel Segmentation

### v1

@@ 3.3 Datasets
U-Net++ 82.653 80.995 97.040 98.450 82.402 79.522 7.6
U-Net3+ Deep 83.006 81.445 97.124 98.427 82.436 79.997 4.1
ResU-Net 82.908 81.299 97.098 98.415 82.481 79.842 6.8
ResU-Net++ 77.008 73.096 96.032 95.312 70.587 71.251 12.8
SAU-Net 82.015 80.170 96.892 98.539 82.195 78.624 9.5
DCASU-Net 82.942 71.335 97.109 98.499 82.383 79.884 7.0
AG-Net 82.249 80.474 96.961 98.566 81.810 78.951 8.6
AttU-Net 83.017 81.448 97.113 98.558 82.830 79.996 2.6
R2U-Net 80.998 78.699 96.808 96.976 77.231 77.122 11.8
ConvU-NeXt 83.015 81.423 97.121 98.525 82.520 79.974 3.8
FR-UNet 82.431 80.709 97.011 98.294 81.677 79.218 9.5
FSG-Net-L (ours) 83.088 81.567 97.106 98.744 83.616 80.121 1.6
Figure 5: Qualitative evaluation on top-3 model structure on DRIVE validation set.
Figure 6: Inspection on output map at each stage $S_{i}$ on FSG-Net. The finer scale of map describes more detailed lines.
3.4 Results
@@
In binary task evaluations, the Matthew correlation coefficient (MCC) is a powerful metric, as noted by Chicco et al. [4]. However, to avoid evaluations oriented towards a specific metric, we also report the average rank of each model, denoted as "Rank Avg" in Table 2. This average rank provides a measure of the stable performance of a model across different datasets. For example, FSG-Net, U-Net3+ [13], and AttU-Net [17] achieved high ranks in all three datasets, whereas ResU-Net and FR-UNet recorded inconsistent results across the three datasets. FSG-Net achieved the highest rank in terms of performance across all three datasets, recording dominant scores in mIoU, F1 score, and MCC, which is equivalent to a detailed expression of the segmentation map. Notably, FSG-Net outperformed the DRIVE dataset, recording SOTA in the F1 score and sensitivity. Figure 5 represents the predicted segmentation maps of the three best models in our evaluation metrics. The FSG-Net shows the best results, especially in segmenting thin vessels. FSG-Net incorporates deep supervision, which allows for the extraction of segmentation maps at each stage. Figure 6 shows the vessel detection results at various stages. It can be observed that as we move to lower levels, the performance in detecting thinner vessels improves. An interesting point is that even at coarse stage $S_{3}$, the detection of thicker vessels is generally well-maintained.
@@
3.5 Ablation study
To further understand the impact of the model capacity and structure on FSG-Net, we conducted ablation studies. In Table 3, we vary the depth of the down-convolution, base channel (Base_c) and structure. The F1 score of three datasets is used as a metric here. The results showed that even the smallest version of FSG-Net surpassed the other models. By comparing the scores in Table 2, the smallest FSG-Net with 1.17M parameters outperformed recent studies with an average rank of 5.2 across all metrics on the three datasets, compared to SAU-Net’s rank of 9.0 with a parameter size of 0.5M, DCASU-Net’s [24] rank of 7.0 with a parameter size of 2.6M, ConvU-NeXt’s rank of 6.7 with a parameter size of 3.5M, and FR-UNet’s rank of 6.6 with a parameter size of 7.4M.
Table 3: Ablation studies on model capacity and model structure
  model Base_c Depth Params (M) DRIVE STARE CHASE_DB1
FSG-Net-L 64 [3, 3, 9, 3] 18.32 83.229 85.100 81.019
@@ 3.5 Ablation study
       Structure
kernel size 3 to 7 35.77 83.006 84.200 80.044
AGF to AttU-Net like 19.61 83.115 84.671 79.333
down-conv to residual 45.60 82.893 82.734 75.765
w/o deep supervision 18.32 83.131 84.698 79.889
removing spatial attention 18.32 83.237 84.598 79.946
To optimize the network structure, we investigated the impact of increasing the kernel size, replacing the attention-guided filter with a simple attention block, and the effect of using pure residual blocks or removing the spatial attention. Our findings show that while increasing the kernel size expands the receptive field, it fails to improve performance in retinal vessel segmentation tasks that require detailed expression of the segmentation map. We found that the attention-guided filter proved to be robust and effective for this task. Additionally, we observed that using pure residual blocks in the deep bottleneck stage increases the parameter size and reduces computational efficiency without significantly improving feature extraction. Removing the spatial attention improved the performance in the DRIVE dataset; however, as the input size increased, the performance decreased proportionally.
@@ 4 Discussion
We previously emphasized the method of feature merging within the network architecture. This is because feature merging allows information from all stages to be mixed and provided as input to the guided filtering module. This approach can be compared to AG-Net, which also gathers features and provides them to the guided filter. Figure 7(a) illustrates the feature merging in the proposed method, while Figure 7(b) shows the corresponding part of AG-Net. Let us focus on the left path to the guided filtering modules in both cases. As can be seen in Figure 7(a), feature maps from three stages (current, upward, and downward) are concatenated and passed through the proposed convolution block. For AG-Net, only the features from the current and lower stages are merged for the attention-guided filter. In summary, our GRM gets feature information from all stages with a modernized convolution, while the AG-Net gets it from current and lower stages with conventional convolutions. For this reason, we were able to achieve superior performance with one fewer stage than AG-Net.
@@
The next discussion point concerns the experimental datasets. DRIVE, STARE, and Chase_DB1 are representative datasets for retinal vessel segmentation. Models that improve the architecture of layers and attention mechanisms continue to emerge, but the number of images in such datasets is limited. Although this paper focuses on retinal vessel segmentation, it is necessary to apply the proposed model to large databases from other fields to verify its variability and scalability. In our future work, we can consider expanding our application areas to include crack detection, facial wrinkle detection, and general medical imaging.
5 Conclusion
@@
In this study, we presented a full-scale representation guided network for retinal vessel segmentation that achieves SOTA on the DRIVE dataset. A modern convolutional block tailored for retinal vessel segmentation was designed, and the guided convolution block maximized performance using full-scale information from the feature representation network. The proposed guided convolution block is compatible with any U-Net architecture, offering scalability for similar tasks across various domains. We encourage future research to use the FSG-Net architecture and our experimental settings to ensure reproducibility and scalability in similar tasks.

### v2

@@ Introduction
U-Net++ 82.653 80.995 97.040 98.450 82.402 79.522 7.8
U-Net3+ Deep 83.006 81.445 97.124 98.427 82.436 79.997 4.4
ResU-Net 82.908 81.299 97.098 98.415 82.481 79.842 7.0
ResU-Net++ 77.008 73.096 96.032 95.312 70.587 71.251 13.8
SAU-Net 82.015 80.170 96.892 98.539 82.195 78.624 10.3
DCASU-Net 82.942 71.335 97.109 98.499 82.383 79.884 7.3
AG-Net 82.249 80.474 96.961 98.566 81.810 78.951 8.7
AttU-Net 83.017 81.448 97.113 98.558 82.830 79.996 2.8
R2U-Net 80.998 78.699 96.808 96.976 77.231 77.122 12.8
ConvU-NeXt 83.015 81.423 97.121 98.525 82.520 79.974 4.0
FR-UNet 82.431 80.709 97.011 98.294 81.677 79.218 9.7
HRNet 82.067 80.226 96.944 98.563 80.957 78.668 9.8
FSG-Net (ours) 83.088 81.567 97.106 98.744 83.616 80.121 1.6
Results
@@
In binary task evaluations, the Matthew correlation coefficient (MCC) is a powerful metric, as noted by Chicco et al. [Chicco2021TheBO]. However, to avoid evaluations oriented towards a specific metric, we also report the average rank of each model, denoted as "Rank Avg" in Table 2. This average rank provides a measure of the stable performance of a model across different datasets. For example, FSG-Net, U-Net3+ [9053405], and AttU-Net [DBLP:journals/corr/abs-1804-03999] achieved high ranks in all four datasets, whereas ResU-Net, FR-UNet and HRNet [Sun_2019_CVPR] recorded inconsistent results across the three datasets. FSG-Net consistently demonstrated top-tier performance across all four datasets, recording dominant scores in mIoU, F1 score, and MCC, which is equivalent to a detailed expression of the segmentation map. Notably, FSG-Net outperformed previous methods on the DRIVE dataset, achieving SOTA performance in F1 score and sensitivity.
We conducted comparative experiments with existing architectures that aim to preserve full-resolution features, namely HRNet and FR-UNet. As shown in Table 2, HRNet achieves the best performance on the CHASE_DB1 dataset and also records the highest accuracy on the DRIVE and STARE datasets. However, its performance noticeably drops on the HRF dataset, and notably, HRNet contains 65.81M parameters, which significantly exceeds the model capacities of all other compared architectures that maintain parameter counts below 30M. FR-UNet performs well on the DRIVE dataset but shows relatively moderate results on the remaining datasets. These comparisons with high-resolution preserving models demonstrate the performance stability and robustness of FSG-Net across diverse datasets, validating the effectiveness of utilizing full-scale information at multiple decoding stages.
Table 3: Performance comparison of ViT-based models and FSG-Net on four retinal vessel segmentation datasets.
Architecture
mIoU
F1 score
Acc
AUC
Sen
MCC
DRIVE
Swin 78.758 76.040 96.296 97.203 75.681 74.171
MaxViT-512 79.702 77.738 96.050 98.452 79.097 75.693
FSG-Net (ours) 18.32 84.068 83.229 97.042 98.235 84.207 81.731
STARE
Swin 79.354 75.900 96.958 97.759 75.186 74.356
MaxViT-512 81.072 78.898 96.816 97.865 79.667 77.357
FSG-Net (ours) 86.118 85.100 97.746 98.967 86.608 83.958
CHASE_DB1
Swin 77.708 73.850 96.966 99.326 76.569 72.347
MaxViT-512 79.322 76.405 96.978 99.494 79.878 74.909
FSG-Net (ours) 82.680 81.019 97.515 99.378 85.995 79.899
HRF
Swin 78.739 74.771 97.420 98.022 74.831 73.548
MaxViT-512 74.980 70.617 95.346 98.645 72.600 68.204
FSG-Net (ours) 83.088 81.567 97.106 98.744 83.616 80.121
We have also conducted experiments to assess the suitability of ViT-based models for the task of retinal vessel segmentation. As noted in the introduction, the thin and elongated structure of retinal vessels poses specific challenges, particularly for models that rely heavily on global representations such as pure ViT. To explore this further, we evaluated Swin-T [Liu_2021_ICCV], which introduces a hierarchical representation with shifted window attention, and MaxViT-T-512 [10.1007/978-3-031-20053-3_27], a hybrid model that combines convolution and attention mechanisms and was evaluated using the UPerNet [Xiao_2018_ECCV] decoding head. As shown in Table 3, the ViT-based models in our setting yielded lower overall performance compared to FSG-Net, particularly in metrics such as F1 score and sensitivity. Swin-T and MaxViT-T-512 contain 58.91M and 59.60M parameters, respectively, yet FSG-Net achieves substantially better performance despite operating with a more parameter-efficient design.
Nonetheless, recent advances in DETR-like architectures [10.1007/978-3-030-58452-8_13] have introduced object query mechanisms that demonstrate strong potential for enhancing small object detection. These efforts reflect ongoing attempts to overcome the limitations traditionally associated with representing fine, low-saliency structures using transformer-based models. Although not included in our experiments, TCDDU-Net [lv2024tcddu], which combines a dual-path U-Net architecture with a Swin backbone, has been reported to outperform several CNN-based methods in terms of quantitative metrics and to effectively segment peripheral vessels. Although not directly comparable due to different experimental settings, it is notable that TCDDU-Net achieved an F1 score of 82.65 on the DRIVE dataset, which is slightly lower than the 84.068 recorded by our proposed FSG-Net. These research findings suggest that future research may benefit from further exploring hybrid architectures or improving transformer-based designs to better handle fine-grained and low-saliency structures such as retinal vessels.
Fig. 5 represents the predicted segmentation maps of the three best models in our evaluation metrics. The FSG-Net shows the best results, especially in segmenting thin vessels. Fig. 6 shows prediction results of FSG-Net obtained through deep supervision from the decoder stages indicated in Fig. 2. As shown in Fig. 6, intermediate predictions may not directly provide fine-grained details, but they can improve the final segmentation results due to their accurate semantic-level information and the capability to capture global context. The quantitative performance benefits of such deep supervision is further discussed in the ablation study.
Figure 5: Qualitative comparison of the top-3 performing models on the DRIVE validation set.
Figure 6: Visualization of predictions at each stage $S_{i}$ of FSG-Net. Here, $S_{i}$ represents each decoder stage, and a larger value of $i$ indicates a deeper decoder as shown in Fig. 2. The first four rows show the entire prediction results from various datasets, while the last three rows present magnified views of these predictions.
@@
Table 4: Cross-domain evaluation across datasets, DRIVE (D), CHASE_DB1 (C), STARE (S), HRF (H). $D\rightarrow C$ indicates that the model was trained on the (D) dataset and evaluated on the (C) dataset.
Dataset
mIoU
F1 score
Acc
AUC
Sen
MCC
$D\rightarrow C$ 79.76 (-2.92) 77.31 (-3.70) 96.46 (-1.05) 99.01 (-0.36) 87.84 (+1.85) 76.14 (-3.75)
$D\rightarrow S$ 82.14 (-3.97) 80.29 (-4.81) 96.95 (-0.79) 97.77 (-1.19) 82.93 (-3.67) 78.90 (-5.05)
$D\rightarrow H$ 79.24 (-3.84) 76.45 (-5.11) 96.21 (-0.89) 98.92 (+0.18) 81.93 (-1.68) 74.75 (-5.37)
$C\rightarrow D$ 78.38 (-5.68) 75.56 (-7.66) 96.25 (-0.79) 88.56 (-9.67) 66.98 (-17.2) 74.44 (-7.29)
$C\rightarrow S$ 81.51 (-4.60) 79.54 (-5.56) 96.98 (-0.76) 96.34 (-2.62) 77.52 (-9.08) 78.09 (-5.86)
$C\rightarrow H$ 80.25 (-2.83) 77.77 (-3.79) 96.57 (-0.53) 97.96 (-0.78) 79.86 (-3.75) 76.13 (-3.99)
$S\rightarrow D$ 80.70 (-3.36) 78.82 (-4.40) 96.52 (-0.52) 94.21 (-4.02) 74.84 (-9.36) 77.22 (-4.51)
$S\rightarrow C$ 79.87 (-2.81) 77.50 (-3.51) 96.56 (-0.95) 99.20 (-0.17) 86.04 (+0.05) 76.16 (-3.73)
$S\rightarrow H$ 78.38 (-4.70) 76.01 (-5.55) 96.08 (-1.02) 98.79 (+0.05) 82.49 (-1.12) 74.33 (-5.79)
$H\rightarrow D$ 79.13 (-4.93) 76.64 (-6.58) 96.27 (-0.77) 92.27 (-5.96) 70.79 (-13.4) 75.07 (-6.66)
$H\rightarrow C$ 81.80 (-0.88) 79.88 (-1.13) 97.10 (-0.41) 99.59 (+0.22) 83.69 (-2.30) 78.46 (-1.43)
$H\rightarrow S$ 82.18 (-3.93) 80.15 (-4.95) 97.15 (-0.59) 95.06 (-3.90) 76.81 (-9.79) 78.89 (-5.06)
To investigate the domain generalization capability of our model in retinal vessel segmentation, we conducted cross-domain validation experiments as presented in Table 4. Specifically, we trained the model on the DRIVE dataset and evaluated its performance on different unseen datasets, namely CHASE_DB1 (D $\rightarrow$ C), STARE (D $\rightarrow$ S), and HRF (D $\rightarrow$ H). As expected, performance decreased across most evaluation metrics when the model was tested on datasets different from the training domain. For example, in the D→C scenario, we observed reductions in mIoU (-2.92), F1 score (-3.70), and MCC (-3.75), although sensitivity showed a slight improvement (+1.85). Similar observations were made in other domain pairs, notably D→S and D→H, reflecting the inherent challenges posed by domain shifts such as differences in imaging modalities, resolutions, and contrast distributions.
However, despite this performance drop, the model still maintained reasonable predictive capability, indicating its robustness and capacity to learn transferable, domain-invariant features. These findings underscore the importance of enhancing domain generalization by exploring additional techniques, such as domain augmentation, pre-training strategies, linear probing, or knowledge distillation, to further improve the model’s adaptability to unseen domains.
Ablation study
Table 5: Ablation study on model capacity.
model
Base_c
Depth
Params(M)
DRIVE
STARE
CHASE_DB1
FSG-Net-L (proposed) 64 [3, 3, 9, 3] 18.32 83.229 85.100 81.019
@@
To further understand the impact of the model capacity and structure on FSG-Net, we conducted ablation studies. In Table 5, we vary the depth of the down-convolution, base channel (Base_c) and structure. The F1 score of three datasets is used as a metric here. The results showed that even the FSG-Net-N surpassed the other models. By comparing the scores in Table 2, the FSG-Net-N with parameter size of 1.17M outperformed recent studies with an average rank (5.2) across all metrics on the three datasets, compared to SAU-Net’s rank (10.8) with a parameter size of 0.5M, DCASU-Net’s [https://doi.org/10.48550/arxiv.2202.00972] rank (8.0) with a parameter size of 2.6M, ConvU-NeXt’s [HAN2022109512] rank (6.8) with a parameter size of 3.5M, and FR-UNet’s rank (8.2) with a parameter size of 7.4M.
Table 6: Ablation study on the proposed modules. The abbreviations represent: DC – replacement of the residual block with a down-convolution structure; GRM – integration of the GRM block; SA – spatial attention at the bottleneck; DKS – dynamic kernel size evaluation; DS – use of deep supervision.
  DC GRM SA DKS DS Params(M) DRIVE STARE CHASE_DB1
39.09 82.580 81.786 78.944
16.85 82.254 83.323 79.409
18.31 83.115 84.671 79.333
18.32 83.131 84.698 79.889
35.77 83.006 84.200 80.044
18.32 83.229 85.100 81.019
To validate the contribution of each proposed component, we conducted a series of ablation experiments on the DRIVE, STARE, and CHASE_DB1 datasets. Table 6 presents the ablation study conducted on the main modules of FSG-Net, with each module denoted by its corresponding abbreviation. Starting from the R2U-Net [https://doi.org/10.48550/arxiv.1802.06955] baseline, chosen for its strong generalization across medical image segmentation tasks, we incrementally integrated the modules proposed in FSG-Net. As shown in the second row of Table 6, replacing the residual blocks in R2U-Net with our proposed down-convolution (as illustrated in Fig. 3(c)) led to a slight performance drop on the DRIVE dataset, while improving the results on STARE and CHASE_DB1. However, this modification reduced the number of parameters by more than half, indicating a favorable trade-off between efficiency and accuracy. As can be seen in the third row of the Table 6, substituting the standard U-Net-style feature concatenation with the proposed GRM block resulted in notable performance improvements on both DRIVE and STARE. This confirms that our GRM module enhances multi-scale feature aggregation and representation learning. Introducing a lightweight spatial attention mechanism at the bottleneck stage yielded consistent performance gains with only a marginal increase in parameter count. This demonstrates the module’s effectiveness in enhancing contextual understanding without significant overhead. To assess the impact of kernel size, we compared a series of three 3×3 convolutions with a single 7×7 convolution. While the 7×7 configuration exhibited minor improvements in high-resolution scenarios due to its direct abstraction style, it generally led to increased parameter counts and lower performance on other datasets, suggesting limited generalizability. Incorporating deep supervision, as shown in the sixth row of the Table 6, consistently improved performance, supporting its role in guiding the learning process through additional gradient signals during training.
Discussion
@@
We previously emphasized the method of feature merging within the network architecture. This is because feature merging allows information from all stages to be mixed and provided as input to the guided filtering module. This approach can be compared to AG-Net, which also gathers features and provides them to the guided filter. Fig. 7(a) illustrates the feature merging in the proposed method, while Fig 7(b) shows the corresponding part of AG-Net. Let us focus on the left path to the guided filtering modules in both cases. As can be seen in Fig 7(a), feature maps from three stages (current, upward, and downward) are concatenated and passed through the proposed convolution block. For AG-Net, only the features from the current and lower stages are merged for the attention-guided filter. In summary, our GRM gets feature information from all stages with a modernized convolution, while the AG-Net gets it from current and lower stages with conventional convolutions. For this reason, we were able to achieve superior performance with one fewer stage than AG-Net.
@@
The next discussion point concerns the experimental datasets. DRIVE, STARE, and Chase_DB1 are representative datasets for retinal vessel segmentation. Models that improve the architecture of layers and attention mechanisms continue to emerge, but the number of images in such datasets is limited. Although this paper focuses on retinal vessel segmentation, it is necessary to apply the proposed model to large databases from other fields to verify its variability and scalability. To this end, we plan to extend FSG-Net to a broader spectrum of medical-imaging tasks—including semi-supervised histological image segmentation [jin2022semi, jin2024inter], skin-lesion diagnosis and segmentation [jin2021cascade], and COVID-19 infection delineation in CT images [jin2021domain]—so as to validate its robustness under limited annotations and pronounced domain shifts.
Finally, we discuss recent approaches to retinal vessel segmentation that differ from ours in terms of network architecture and utilization of attention mechanisms. Unlike our proposed FSG-Net, which utilizes outputs inferred from all layers of a U-Net architecture, Tan et al. [tan2024deep] proposed a W-shaped RCAR-UNet structure that aggregates outputs from the preceding U-Net and passes them forward to the next U-Net module. Meanwhile, Ding et al. [ding2024rcar] incorporated a rough set-based channel-attention mechanism into the U-Net structure to reinforce the long-range dependency of retinal vessels. This differs from our approach, which combines information from all scales and passes it through a trainable guided filter. In future studies, we plan to explore integrating these complementary attention and filtering strategies with our full-scale guided network framework, potentially yielding further improvements in segmentation performance.
Conclusion
@@
In this study, we presented a full-scale representation guided network (FSG-Net) for retinal vessel segmentation that demonstrates competitive performance approaching current SOTA methods on the DRIVE dataset. A modern convolutional block tailored for retinal vessel segmentation was designed, and the guided convolution block effectively leveraged full-scale information from the feature representation network. The proposed guided convolution block is compatible with any U-Net architecture, offering scalability for similar tasks across various domains. However, our method remains purely CNN-based, which might limit its capability compared to recent transformer-integrated U-Net architectures designed for retinal vessel segmentation tasks. Therefore, future research should explore integrating transformer modules into our full-scale feature representation approach. We encourage future research to use the FSG-Net architecture and our experimental settings to ensure reproducibility and scalability in similar tasks.